# Aligning Beam with Imbalanced Multi-modality: A Generative Federated Learning Approach

### v2

@@ Abstract
As vehicle intelligence advances, multi-modal sensing-aided communication emerges as a key enabler for reliable Vehicle-to-Everything (V2X) connectivity through precise environmental characterization. As centralized learning may suffer from data privacy, model heterogeneity and communication overhead issues, federated learning (FL) has been introduced to support V2X. However, the practical deployment of FL faces critical challenges: model performance degradation from label imbalance across vehicles and training instability induced by modality disparities in sensor-equipped agents. To overcome these limitations, we propose a generative FL approach for beam selection (GFL4BS). Our solution features two core innovations: 1) An adaptive zero-shot multi-modal generator coupled with spectral-regularized loss functions to enhance the expressiveness of synthetic data compensating for both label scarcity and missing modalities; 2) A hybrid training paradigm integrating feature fusion with decentralized optimization to ensure training resilience while minimizing communication costs. Experimental evaluations demonstrate significant improvements over baselines achieving 16.2% higher accuracy than the current state-of-the-art under severe label imbalance conditions while maintaining over 70% successful rate even when two agents lack both LiDAR and RGB camera inputs.
@@ I Introduction
In 6G integrated sensing-communications frameworks, sensor data enables detailed environmental characterization. Hence, sensing-aided methods guide efficient beam alignment. For example, deep learning techniques leveraging RF data have been explored in [6, 7]. With advancing vehicle intelligence, multi-modal sensors (GPS/LiDAR/RGB cameras) now supplant RF-only solutions due to richer environmental insights [8, 9]. While federated learning (FL) addresses privacy concerns inherent to heterogeneous datasets [10, 11], existing FL strategies neglect modality imbalance caused by diverse sensor configurations (e.g., type/performance), leading to incomplete multi-modal inputs and heightened training instability risks. [12] employs a modality‑agnostic approach to accommodate any combination of sensor inputs while [13] adopts vertical FL to capture cross‑modal feature heterogeneity. Nevertheless, both methods fail to address partial data loss or inter‑vehicle label imbalance.
@@
Specifically, the BS distributes model branches according to vehicle sensor configurations. Each vehicle updates its model branch using its own sensor data, and then sends them to the BS, which aggregates each branch separately to construct a new global model for the next iteration. Simulations indicate that under label imbalance conditions, GFL4BS enhances Top-1 accuracy compared to benchmarks, closely matching the centralized learning method while maintaining high communication performance. Under modality imbalance conditions, GFL4BS not only achieves high accuracy in model training but also reduces communication overhead during training.
@@ II System Model
As shown in Fig. 1(a), we consider a multi-user MIMO system where the BS, equipped with a uniform array comprising $N_{t}$ antennas, serves $V$ single-antenna vehicles. A predefined beamforming codebook $C=\{c^{1},c^{2},\ldots,c^{M}\}$ is employed for beam alignment. The BS uses the beamforming vector $\mathbf{w}_{c^{m},v}\in\mathbb{C}^{N_{t}\times 1}$ corresponding to beam $c^{m}$ to transmit the normalized data symbol $s_{v}$ to the $v$-th vehicle, subject to the power constraint $\sum_{v=1}^{V}\|\mathbf{w}_{c^{m},v}\|^{2}\leq P$. The received signal at the $v$-th vehicle is expressed as:
@@
The modality processing branches shown in Fig. 1(b) are specifically tailored to the unique characteristics of each modality. Let $N_{v}$ denote the number of samples for each vehicle. Then, the inputs are defined as follows: the GPS data $\mathbf{x}_{v}^{\text{GPS}}\in\mathbb{R}^{N_{v}\times 2}$, the RGB data $\mathbf{x}_{v}^{\text{RGB}}\in\mathbb{R}^{N_{v}\times d_{0}^{\text{R}}\times d_{1}^{\text{R}}}$ and the LiDAR data $\mathbf{x}_{v}^{\text{L}}\in\mathbb{R}^{N_{v}\times d_{0}^{\text{L}}\times d_{1}^{\text{L}}\times d_{2}^{\text{L}}}$. Accordingly, the GPS data provides 2D coordinates representing the relative distance between the vehicle and the BS, thereby conveying Line-of-Sight (LoS) information; the RGB data contains rich environmental details, including reflectors and obstacles that contribute to multipath; and the LiDAR data, which is preprocessed into cuboidal regions based on its distance from the BS [^1] Each cuboidal region is assigned a unique label: blocking obstacles are denoted by $1$, the transmitter by $-1$, and the receiver by $-2$[11]. depicts the spatial structural information of the communication environment. To maintain model stability during both training and inference, every conventional and fully connected layer is followed by a batch normalization (BN) layer. The uni-modal processing branches extract modality features $\mathbf{f}^{\text{G}}$, $\mathbf{f}^{\text{R}}$, and $\mathbf{f}^{\text{L}}$. These features are subsequently concatenated to capture correlations between the raw data and the beam index.
@@ III-A Adaptive Zero-shot Multi-modal Data Generator
To preserve data privacy, it is impractical to train a centralized generator that requires access to data to determine the global distribution in advance. Inspired by [17], we propose AMD generator that can fulfill multi-modal data generation requirements. During the model training period, each BN layer normalizes the forward activations by computing the mini-batch mean and variance to reduce internal covariate shift and accelerate training, while updating the BN-specific parameters, namely the running mean $\mu$ and running variance $\sigma$. These parameters reflect the distribution information of the input data. Therefore, the goal of data generation can be described as finding synthetic data ${\hat{\mathbb{D}}}=\{\hat{\mathrm{d}}^{1},\hat{\mathrm{d}}^{2},\ldots,\hat{\mathrm{d}}^{N}\}$ that results in similar statistics to those in the BN layers of the global model. Here, $N$ represents the number of data samples and $\hat{\mathrm{d}}^{n}$ is a tuple consisting $[\hat{x}^{\text{G},n},\hat{x}^{\text{R},n},\hat{x}^{\text{L},n}]$. The statistic $\hat{\mu}_{l}$ and $\hat{\sigma}_{l}$ of the activation $a_{l}$ in the forward pass of the $l$-th layer are calculated by:
@@
$\hat{\mu}_{l}(\hat{\mathbb{D}})$ $=\frac{1}{N}\sum_{\hat{\mathrm{d}}^{n}\in\hat{\mathbb{D}}}a_{l}(\hat{\mathrm{d}}^{n}),$ (3a)
$\hat{\sigma}_{l}(\hat{\mathbb{D}})$ $=\frac{1}{N}\sum_{\hat{\mathrm{d}}^{n}\in\hat{\mathbb{D}}}\left(a_{l}(\hat{\mathrm{d}}^{n})-\hat{\mu}_{l}(\hat{\mathbb{D}})\right)^{2}.$ (3b)
@@
Here, $\hat{\mathbf{c}}_{p}=\mathcal{M}(\hat{\mathbb{D}})$ represents the prediction of the global model. Gradient descent is employed to solve Eq. (5). Specifically, the synthetic input is randomly initialized from a Gaussian distribution and iteratively updated to minimize $\mathcal{L}_{gen}$. Furthermore, Eq. (5) can be adjusted to fulfill the requirement for modality data filling, involving three steps. First, based on the modality type, the initialization is modified from $[\hat{x}^{\text{G},n},\hat{x}^{\text{R},n},\hat{x}^{\text{L},n}]$ to a combination of existing raw data and synthetic features $[{x}^{\text{Existed},n},\hat{f}^{\text{Missing},n}]$. Second, the $L^{bn}$ BN layers in the global model are replaced with $L^{bn}_{\text{I}}$ BS layers in the integration branch. These two adjustments aim to reduce discrepancies in BN statistics among various extractor branches. Finally, the hard label constraint $\mathcal{H}(\hat{\mathbf{c}}_{\text{hard}},\hat{\mathbf{c}}_{p})$ is replaced with true label constraint $\mathcal{H}(\hat{\mathbf{c}},\hat{\mathbf{c}}_{p})$ because the label $\mathbf{c}$ is known in data filling task.
@@
By doing so, after each update, we embed the receiver and transmitter positions as 22–2– 2 and 11–1– 1, respectively, ensuring that the synthetic data format matches the original.
@@ III-B Quantification of Label and Modality Imbalances
Accordingly, each vehicle augments their data with synthetic data generated by AMD generator to drive $\zeta$, $\epsilon$ and $\kappa_{v}^{\text{Q}}$ closer to one. This approach helps mitigate model performance degradation and reduces the risk of training failure by balancing the data distribution across different modalities and vehicles.
@@ III-C Distributed Training Scheme
[…]
@@
Figure 3: Results of GFL4BS in different level modality imbalance.
@@ IV-A Experiment Setting
1) Label-Imbalanced Dataset: The original dataset exhibits the label imbalance, with average overlap rate: $\zeta=\textbf{78.9\%}$%. To further evaluate GFL, there are three datasets with varying degrees of imbalance, namely Low (L), Medium (M), and High (H). The labeled data are first ranked by volume and then split into two equal groups (top 50% and bottom 50%). One of these groups is randomly selected for data removal based on a 7:3 probability ratio. Specifically, the numbers of data points removed for the L, M, and H datasets are 6, 9 and 12, respectively. These labeled data are transferred to the next client until the final client receives the additional data. This strategy not only increases the imbalance level but also mitigates any negative effects from reducing dataset size. To minimize the influence of random label selection, the set of removed labels in the higher-imbalance datasets is constructed based on those of the lower-imbalance dataset. The $\zeta$ values for the datasets are 44.3%, 35.1%, and 24.7%.
@@ IV-E Communication Performance
Fig. 4 illustrates communication performance results. GFL4BS yields a 96.8% sum rate ratio at an average overlap rate of $\zeta$=78.9and a 92.9% sum rate ratio when 80% of the RGB and LiDAR data are missing. Communication overhead during training is quantified by defining an overhead ratio relative to communication performance of FLASH. As shown, under label imbalance, GFL4BS requires fewer resources because additional data reduces the number of communication rounds. FLASH does not exhibit an advantage, since fewer local training epochs result in slower convergence. The two vehicles lacking LiDAR and RGB cameras transmit the GPS extractor and integration branches to the BS, which account for approximately 25% of the model. Consequently, GFL4BS reduces communication overhead per training round by 15%.

### v3

@@ Abstract
As vehicle intelligence advances, multi-modal sensing-aided communication emerges as a key enabler for reliable Vehicle-to-Everything (V2X) connectivity through precise environmental characterization. As centralized learning may suffer from data privacy, model heterogeneity and communication overhead issues, federated learning (FL) has been introduced to support V2X. However, the practical deployment of FL faces critical challenges: model performance degradation from label imbalance across vehicles and training instability induced by modality disparities in sensor-equipped agents. To overcome these limitations, we propose a generative FL approach for beam selection (GFL4BS). Our solution features two core innovations: 1) An adaptive zero-shot multi-modal generator coupled with spectral-regularized loss functions to enhance the expressiveness of synthetic data compensating for both label scarcity and missing modalities; 2) A hybrid training paradigm integrating feature fusion with decentralized optimization to ensure training resilience while minimizing communication costs. Experimental evaluations demonstrate significant improvements over baselines achieving 16.2% higher accuracy than the current state-of-the-art under severe label imbalance conditions while maintaining over 70% successful rate even when two agents lack both LiDAR and RGB camera inputs. Our code is available at the https://github.com/Jiahui-L/GFL4BS.
@@ I Introduction
In 6G integrated sensing-communications frameworks, sensor data enables detailed environmental characterization. Hence, sensing-aided methods guide efficient beam alignment. For example, deep learning techniques leveraging RF data have been explored in [6, 7, 8]. With advancing vehicle intelligence, multi-modal sensors now supplant RF-only solutions due to richer environmental insights [9, 10]. While federated learning (FL) addresses privacy concerns inherent to heterogeneous datasets [11, 12], existing FL strategies neglect modality imbalance caused by diverse sensor configurations (e.g., type/performance), leading to incomplete multi-modal inputs and heightened training instability risks. [13] employs a modality‑agnostic approach to accommodate any combination of sensor inputs while [14] adopts vertical FL to capture cross‑modal feature heterogeneity. Nevertheless, both methods fail to address partial data loss or inter‑vehicle label imbalance.
@@
Specifically, the BS distributes model branches according to vehicle sensor configurations. Each vehicle updates its model branches using its own sensor data, and then sends them to the BS, which aggregates each branch separately to construct a new global model for the next iteration. Simulations indicate that under label imbalance conditions, GFL4BS enhances Top-1 accuracy compared to benchmarks, closely matching the centralized learning method while maintaining high communication performance. Under modality imbalance conditions, GFL4BS not only achieves high accuracy in model training but also reduces communication overhead during training.
@@ II System Model
As shown in Fig. 1(a), we consider a multi-user multiple-input multiple-output (MIMO) system where the BS, equipped with an $N_{t}$ uniform array, serves $V$ single-antenna vehicles. A predefined beamforming codebook $C=\{c^{1},c^{2},\ldots,c^{M}\}$ is employed for beam alignment. The BS uses the beamforming vector $\mathbf{w}_{c^{m},v}\in\mathbb{C}^{N_{t}\times 1}$ corresponding to beam $c^{m}$ to transmit the normalized data symbol $s_{v}$ to the $v$-th vehicle, subject to the power constraint $\sum_{v=1}^{V}\|\mathbf{w}_{c^{m},v}\|^{2}\leq P$. The received signal at the $v$-th vehicle is expressed as:
@@
The modality processing branches shown in Fig. 1(b) are specifically tailored to the unique characteristics of each modality. Let $N_{v}$ denote the number of samples for each vehicle. Then, the inputs are defined as follows: the GPS data $\mathbf{x}_{v}^{\text{GPS}}\in\mathbb{R}^{N_{v}\times 2}$, the RGB data $\mathbf{x}_{v}^{\text{RGB}}\in\mathbb{R}^{N_{v}\times d_{0}^{\text{R}}\times d_{1}^{\text{R}}}$ and the LiDAR data $\mathbf{x}_{v}^{\text{L}}\in\mathbb{R}^{N_{v}\times d_{0}^{\text{L}}\times d_{1}^{\text{L}}\times d_{2}^{\text{L}}}$. Accordingly, the GPS data provides 2D coordinates representing the relative distance between the vehicle and the BS, thereby conveying Line-of-Sight (LoS) information; the RGB data contains rich environmental details, including reflectors and obstacles that contribute to multipath; and the LiDAR data, which is preprocessed into cuboidal regions based on its distance from the BS [^1] Each cuboidal region is assigned a unique label: blocking obstacles are denoted by $1$, the transmitter by 1-1, and the receiver by 2-2[12]., depicts the spatial structural information of the communication environment. To maintain model stability during both training and inference, every conventional and fully connected layer is followed by a batch normalization (BN) layer. The uni-modal processing branches extract modality features $\mathbf{f}^{\text{G}}$, $\mathbf{f}^{\text{R}}$, and $\mathbf{f}^{\text{L}}$. These features are subsequently concatenated to capture correlations between the raw data and the beam index.
@@ III-A Adaptive Zero-shot Multi-modal Data Generator
To preserve data privacy, it is impractical to train a centralized generator that requires access to data to determine the global distribution in advance. Inspired by [18], we propose AMD generator that can fulfill multi-modal data generation requirements. During the model training period, each BN layer normalizes the forward activations by computing the mini-batch mean and variance to reduce internal covariate shift and accelerate training, while updating the BN-specific parameters, namely the running mean $\mu$ and running variance $\sigma$. These parameters reflect the distribution information of the input data [19]. Therefore, the goal of data generation can be described as finding synthetic data ${\hat{\mathbb{D}}}=\{\hat{\mathrm{d}}^{1},\hat{\mathrm{d}}^{2},\ldots,\hat{\mathrm{d}}^{N}\}$ that results in similar statistics to those in the BN layers of the global model. Here, $N$ represents the number of samples and $\hat{\mathrm{d}}^{n}$ is a tuple consisting $\hat{x}^{\text{G},n},\hat{x}^{\text{R},n},\hat{x}^{\text{L},n}]$. The $\hat{\mu}_{l}$ and $\hat{\sigma}_{l}$ of the activation $a_{l}$ in the forward pass of the $l$-th layer are calculated by:
@@
[…]
@@
[…]
@@
By doing so, after each update, we embed the receiver and transmitter positions as 2–2 and 1–1, respectively, ensuring that the synthetic data format matches the original.
@@ III-B Quantification of Label and Modality Imbalances
Accordingly, each vehicle augments its data with synthetic data generated by AMD generator to drive $\zeta$, $\epsilon$ and $\kappa_{v}^{\text{Q}}$ closer to one. This approach helps mitigate model performance degradation and reduces the risk of training failure by balancing the data distribution across different modalities and vehicles.
@@ III-C Distributed Training Scheme
$\text{Integration:}\ \mathbf{\Theta}^{\text{I}}_{v,t+1}$ $=\mathbf{\Theta}^{\text{I}}_{v,t}-\eta\,\nabla_{\mathbf{\Theta}^{\text{I}}_{v,t}}\mathcal{L}_{b}\bigl{(}\mathbf{\Theta}^{\text{I}}_{v,t};\,\mathbf{f}^{\text{I}}\bigr{)},$ (11a)
$\text{Extractor:}\ \mathbf{\Theta}^{\text{Q}}_{v,t+1}$ $=\mathbf{\Theta}^{\text{Q}}_{v,t}-\eta\,\nabla_{\mathbf{\Theta}^{\text{Q}}_{v,t}}\mathcal{L}_{b}\bigl{(}\mathbf{\Theta}^{\text{Q}}_{v,t};\,\mathbb{D}_{v}\bigr{)}.$ (11b)
@@
Figure 3: Results of GFL4BS in different level modality imbalance.
@@ IV-A Experiment Setting
1) Label-Imbalanced Dataset: The original dataset exhibits the label imbalance, with average overlap rate: $\zeta=\textbf{78.9\%}$. To further evaluate GFL, there are three datasets with varying degrees of imbalance, namely Low (L), Medium (M), and High (H). The labeled data are first ranked by volume and then split into two equal groups (top 50% and bottom 50%). One of these groups is randomly selected for data removal based on a 7:3 probability ratio. Specifically, the numbers of data points removed for the L, M, and H datasets are 6, 9 and 12, respectively. These labeled data are transferred to the next client until the final client receives the additional data. This strategy not only increases the imbalance level but also mitigates any negative effects from reducing dataset size. To minimize the influence of random label selection, the set of removed labels in the higher-imbalance datasets is constructed based on those of the lower-imbalance dataset. The $\zeta$ values for the datasets are 44.3%, 35.1%, and 24.7%.
@@ IV-E Communication Performance
Fig. 4 illustrates communication performance results. GFL4BS yields a 96.8% sum rate ratio at an average overlap rate of $\zeta$=78.9% and a 92.9% sum rate ratio when 80% of the RGB and LiDAR data are missing. Communication overhead during training is quantified by defining an overhead ratio relative to communication performance of FLASH. As shown, under label imbalance, GFL4BS requires fewer resources because additional data reduces the number of communication rounds. FLASH does not exhibit an advantage, since fewer local training epochs result in slower convergence. The two vehicles lacking LiDAR and RGB cameras transmit the GPS extractor and integration branches to the BS, which account for approximately 25% of the model. Consequently, GFL4BS reduces communication overhead per training round by 15%.